Session: Feb 2011 (session 1)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Mariotti at Tue Mar 01 04:32:53 2011) Success:
Weather: rain. RFI: as usual, some RFI on band

Torun (K. Borkowski at Tue Mar 01 06:35:14 2011) Success:
No known problems.

Westerbork (Geert Kuper at Tue Mar 01 10:18:36 2011) Success:
Apart from RFI no known errors.

Effelsberg (Uwe Bach at Tue Mar 01 14:35:20 2011) Success:

Urumqi (Zhang hua at Wed Mar 02 02:44:06 2011) Success:
No known problems.

Shanghai (Bo Xia at Wed Mar 02 05:01:07 2011) Success:

Onsala (Michael Lindqvist at Wed Mar 02 05:24:24 2011) Success:
No known problems.

Effelsberg (Uwe Bach at Wed Mar 02 14:38:12 2011) Success:
DBBC: Successfully observed, but likely swapped pols. at DBBC. MKIV was correct.

Jodrell_Bank (Alastair Gunn at Tue Mar 29 11:28:37 2011) Minor failures: ANT
Brief loss of data at 03:23 due to control problem. No other known problems.